PROCEEDINGS OF THE SECOND FUJITSU-ANU CAP WORKSHOP

RICHARD P. BRENT (EDITOR)

Abstract

The Proceedings of a Workshop held at the Australian National University in November 1991. The topic was software and applications for the Fujitsu AP1000 MIMD parallel computer (also known as the CAP II).

Comments

Only the Abstract is given here. The full Proceedings appeared as [2]. The workshop was the second in a series; the Proceedings of the first appeared as [1].

References

1. R. P. Brent and M. Ishii (editors), Proceedings of the First Fujitsu-ANU CAP Workshop, Fujitsu Research Laboratories, Kawasaki, Japan, November 1990, 205 pp. rpb123.

2. R. P. Brent (editor), Proceedings of the Second Fujitsu-ANU CAP Workshop, Australian National University, November 1991, 254 pp. rpb129.

Computer Sciences Laboratory, Australian National University, Canberra E-mail address: rpb@cslab.anu.edu.au

1991 Mathematics Subject Classification. Primary 68-06; Secondary 65-06, 65Y05, 65Y10, 68N20, 68Q22.

Key words and phrases. Parallel computation, Fujitsu AP1000, workshop. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, ANU and Fujitsu Laboratories Ltd. Comments \begin{matrix}c \\ ◯ \end{matrix} 1993, R. P. Brent.

rpb129a typeset using AMS-L^{A}TEX